OpenStudy (anonymous):

Solve for x in the proportion 3 over the quantity x plus 2 equals 6 over the quantity x minus 1. x = −5 x = 1 x = 4 x = 5

OpenStudy (anonymous):

first step is to write the equation \[\huge \frac 3{x+2} = \frac 6{x-1}\]

OpenStudy (anonymous):

then cross multiply. what do you get?

OpenStudy (anonymous):

12 and 3

OpenStudy (johnweldon1993):

\[3(x - 1) = ?\] \[6(x + 2) =?\]

OpenStudy (anonymous):

-3 and 12?

OpenStudy (johnweldon1993):

Don't forget about the 'x' part When you multiply 3 times (x - 1) you multiply 3 by x....and then 3 by -1 so \[3(x - 1) = 3x - 3\] Right?

OpenStudy (johnweldon1993):

So use that same logic with \[6(x + 2) = ?\]

OpenStudy (johnweldon1993):

Did that make sense..??

OpenStudy (anonymous):

So 6x + 12

OpenStudy (anonymous):

Yeah but how do i get the answeR?

OpenStudy (johnweldon1993):

Yes @ahoward79 6x + 12 so combining what I showed you and what you got...now you have \[3x - 3 = 6x + 12\] can you solve for 'x' from here?

OpenStudy (anonymous):

Uhh, can you show me :/

OpenStudy (johnweldon1993):

Sure...don't get discouraged! :)...but let me ask you...do you understand everything up until now..?

OpenStudy (johnweldon1993):

I'll break it down just in case Step 1: Write the equation \[\frac{ 3 }{ x + 2 } = \frac{ 6 }{ x - 1 }\] Step 2...when you have 2 fractions like this equal to each other...you cross multiply Multiply \[3(x - 1) = 6(x + 2)\] *see how it crossed? Step 3 Distribute \[3(x - 1) = (3 \times x)+(3 \times -1) = 3x - 3\] and \[6(x + 2) = (6 \times x)+(6 \times 2) = 6x + 12\] so now you have \[3x - 3 = 6x + 12\] Step 4: Combining like terms... you can only combine numbers with an 'x' and numbers that are constants...so lets get all the numbers with an 'x' on 1 side....and the other numbers on the other Lets start by adding 3 to both sides \[3x -3 + 3 = 6x + 12 + 3\] becomes \[3x = 6x + 15\] Now lets subtract 6x from both sides \[3x - 6x = 6x - 6x + 15\] becomes \[-3x = 15\] can you solve for 'x' from here?

OpenStudy (anonymous):

LOL